Adult rattlesnakes tend to hide rather than bite humans. Out of pure curiosity, humans often go in search of rattlesnakes. Bite records seem to show that rattlesnakes are aware that humans are too big to eat. Medical records of rattlesnake bites show that about 30% of the bites contain no venom. These are known as dry bites. Note: Consider rattlesnake bites independent events.

a) What is the probability that two rattlesnake victims get dry bites?

b) If you are bitten by a rattlesnake, what is the probability the bite contains venom?

c) If an emergency center in West Texas sees six patients with rattlesnake bites during the fall rattlesnake round-up, what is the probability the first five bites contain venom and the last one does not contain venom?
Answer #1

a)probability that two rattlesnake victims get dry bites =P(first one was dry bite and second one was dry bite)

=0.3*0.3 =0.09

b) probability the bite contains venom =1-P(dry bite) =1-0.3 =0.70

c)probability the first five bites contain venom and the last one does not contain venom =(0.7)5*0.3 =0.0504